As a designner you have to obtain results using resistors to follow equation :

R3 <- R1 + R2

What would be an appropriate circuit to do so and why?



1
Expert's answer
2021-09-26T11:59:13-0400

The equivalent resistance of the series connection is:

"Req\n\n\u200b=R+R+R+R=4Rohm"


The total current in the circuit ,

"I=V\/R=12\/4R=3\/R"


Reading of voltmeter A = Voltage across R1

"=I\u00d7R1\\\\\n\n\u200b=3\/R\u00d7R=3V"


Reading of voltmeter B = Voltage across R2

"\u200b=I\u00d7R2\\\\\n\n\u200b=3\/R\u00d7R=3V"


Reading of voltmeter C = Voltage across the series combination of R3

​ and R4

"\u200b=I\u00d7(R3\u200b+R4\n\n\u200b)=3\/R\u00d72R=6V"